OpenStudy (anonymous):

Find complex numbers z, w which simultaneously satisfy: (1 – i)z + 2w = 2 2z + (1 + i)w = 1

OpenStudy (anonymous):

Isolate w on the first equation, wich gets (2-(1-i)z)/2, then replace on the second wich gets 2z+(1+i)(2-(1-i)z)/2=1 and z=3/2-i and w=9/4+2i.